# Sample from What You See: Visuomotor Policy Learning via Diffusion Bridge with Observation-Embedded Stochastic Differential Equation

### v1

@@ 3.2 Modality Fusion and Alignment
Encoding the Robot State and Point Cloud. Our encoding pipeline begins by converting depth images into 3D point clouds, chosen for the efficiency. We then downsample them (512 or 1024 points in simulation; 2048 in real-world) using farthest point sampling [qi2017pointnet]. Finally, a lightweight MLP with LayerNorm and an MLP state encoder map the point cloud and robot state into a shared latent representation for fusion, which can be formulated as:
@@
Figure 3: Examples of the simulation tasks.
Multi-Modality Fusion. To enable the diffusion bridge to map between heterogeneous modalities effectively, we introduce the multi-modality fusion module that integrates the 3D point cloud and robot state into a unified representation. This allows the policy to process diverse inputs coherently within our framework. Specifically, we employ the Cross-Attention mechanism [rombach2022high] to fuse the modalities:
$\boldsymbol{a}_{T}:=\boldsymbol{z}_{obs}=\mathbf{softmax}(\frac{\boldsymbol{z}_{s}\boldsymbol{z}_{s}^{T}}{\sqrt{d_{s}}})\boldsymbol{z}_{pc},$ (6)
@@
Figure 4: Examples of the real-world tasks.
Modality Alignment. Although multi-modality fusion aligns the observation and action distributions in shape, significant distributional differences still remain and the learned observation representation lacks sampling capability. To address this, we propose using the contrastive learning loss to train the aligner. The contrastive loss aligns the semantic proximity of the observation and action distributions. Specifically, we adopt CLIP loss [radford2021learning] as the contrastive learning loss:
@@
DP [chi2023diffusion] 0.79±0.27 0.31±0.24 0.10±0.12 0.26±0.25 0.45±0.08 0.31±0.17 0.37
DP3 [ze20243d] 0.87±0.20 0.61±0.29 0.40±0.37 0.51±0.33 0.57±0.08 0.68±0.06 0.60
Simple DP3 [ze20243d] 0.86±0.12 0.59±0.22 0.38±0.26 0.47±0.28 0.48±0.07 0.68±0.05 0.58
FlowPolicy [zhang2025flowpolicy] 0.86±0.16 0.67±0.21 0.59±0.30 0.76±0.26 0.54±0.09 0.70±0.12 0.68
[…]
@@ 4.1 Simulation Experiment
Oven-Closing Oven-Opening Pick Place Pour Unplug Average
@@
Simulation Benchmark. In simulation tasks, we evaluate our BridgePolicy on 52 tasks across three benchmarks including Adroit [rajeswaran2017learning], DexArt [bao2023dexart], and MetaWorld [yu2020meta]. Adroit and DexArt focus on dexterous hand manipulation with varying complexity, while MetaWorld offers a broad spectrum of robotic arm tasks across different difficulty levels. Examples of tasks on these three benchmarks are illustrated in Figure 3.
@@
Expert Demonstration Collection. We collect expert data with script policy in MetaWorld and adopt the VRL3 [wang2022vrl3] and PPO [schulman2017proximal], two reinforcement learning (RL) algorithms, to collect expert data for Adroit and DexArt, respectively. We collect 10 episodes for tasks of the Adroit and the MetaWorld benchmarks and 100 episodes for the DexArt benchmark to train the policies.
@@
Baselines. For comparison, we select state-of-the-art 2D-based methods DP [chi2023diffusion] which takes the 2D image as its visual input and 3D conditional diffusion or flow based approaches including DP3 [ze20243d], simple DP3 (its lightweight version), and FlowPolicy [zhang2025flowpolicy], which generates the action from random Gaussian noise and leverage 3D point cloud as the visual condition in the neural network as baselines. We set the number of function evaluation (NFE) to 10 for all baselines except FlowPolicy. Based on the settings in the original FlowPolicy paper, we set its NFE to 1 instead of 10 to avoid the error accumulation issue that occurs in multi-step sampling of consistency flow matching [sabour2025align]. For further details of the NFE setting, please refer to the Supplementary Materials.
@@
Evaluation and Implementation. We run 3 random seeds for each experiment, followed by DP3 [ze20243d] and FlowPolicy [zhang2025flowpolicy]. For each random seed, we train the models for 3000 epochs and evaluate the success rate on 20 episodes every 200 training epochs. We pick the five highest success rates per seed and report the mean success rate of the 3 random seeds. For fair comparison, we keep the model architecture consistent with the baselines, the amount of parameters comparable, and the shared hyperparameters, such as learning rates, weight decays of the optimizer, batch sizes, and training epochs. For each task, BridgePolicy is trained on a single NVIDIA RTX 4090 GPU. For further details of the implementation of BridgePolicy and other baselines, please refer to the Supplementary Materials.
@@
Performance in Simulator. For Adroit and DexArt, we report the average success rates directly. Particularly for MetaWorld tasks, it is further categorized into four difficulty levels, with average rates reported for each. The quantitative results are reported in Table 1. Our proposed BridgePolicy demonstrates a consistent performance across all evaluated benchmarks, achieving the highest average success rates. All baseline methods show considerably lower success rates, particularly in more challenging settings (Hard and Very Hard in MetaWorld), which also showcases the robustness of our method in tasks with varying complexity. These results validate the effectiveness of BridgePolicy in leveraging the diffusion bridge formulation for improved policy learning in diverse simulation environments. For success rates of individual tasks, please refer to the Supplementary Materials.
@@ 4.2 Real-world Experiment
Experiment Setup and Evaluation. We evaluate the BridgePolicy on 5 real-world tasks using the Franka Emika Panda robot. The point cloud is acquired using the ZED-2i camera. The real-world tasks are shown in Figure 4. We evaluate the policies on 10 episodes of each task. The implementation remains the same as that in the experiments in simulation. We only increase the number of points of the point cloud to 2048 for a more dense representation of the real-world scenario. For experiment environment setup details of the real-world experiments, please refer to the Supplementary Materials.
Expert Demonstration Collection. The expert demonstrations are collected by the GELLO human teleopertation system [wu2024gello], manipulated by an experienced graduate. For each task, we collect 50 episodes to train the policy. Similarly, in the simulator and for a fair comparison, the expert demonstrations provided to all baseline methods and our BridgePolicy are identical. Please refer to the Supplementary Materials for setup details of the real-world data collection.
@@
Qualitive Comparison on Real-World Task. As depicted in Figure 5 and  7, we select three examples of the real-world task Oven-Opening and Unplug generated by DP3, FlowPolicy, and BridgePolicy, respectively. From left to right, the sequence presents the key frames of the Franka arm finishing the task. The Oven-Opening task consists of 2 key stages: 1) grabbing the door handle of the oven and opening it halfway; 2) moving the arm upon the half-opened door and pushing it down until it is fully opened. In the very case of Oven-Opening, DP3 fails to finish the task while FlowPolicy and BridgePolicy both finish the task in the end. BridgePolicy exactly finishes the 2 stages and fully opens the oven, while FlowPolicy fails to half-open the door in the first stage, and it presses down against the door handle directly, leaving the door not fully opened. We also observe similar results in the Unplug task, shown in Figure 7. DP3 and FlowPolicy are able to extend their grippers towards the target location, but they did not precisely clamp the socket, thus failing to unplug successfully. BridgePolicy, on the other hand, successfully clamped the plug and pulled it out. Videos of the real-world experiments can be found in the Supplementary Materials.
@@
0.3 0.6 1.0
Adroit Door 0.81 0.88 0.81
Adroit Pen 0.665 0.615 0.61
MetaWorld Box-Close 0.74 0.75 0.76
MetaWorld Coffee-Pull 0.95 0.86 0.84
MetaWorld Dial-Turn 0.91 0.87 0.85
MetaWorld Handle-Pull 0.63 0.53 0.49
Avg Success Rate 0.78 0.75 0.72
@@ 4.3 Ablation Study
In this section, we analyze the performance of BridgePolicy under different hyperparameter settings. Additional ablation studies are provided in the Supplementary Materials.
@@
Sample Efficiency. The success rate of the agent in accomplishing tasks largely depends on the number of expert demonstrations. Here, we evaluate how the number of expert demonstrations would affect the success rate of the agent finishing the tasks. As shown in Figure 6, we select four tasks for this ablation study: the MetaWorld Hand-Pull-Side, Pick Place, Stick Pull and the Adroit Pen task. For all four tasks, increasing the number of expert demonstrations generally improves the agent’s success rate. BridgePolicy shows the most stable and robust performance across all demonstration counts among the evaluated tasks, especially with fewer training samples. For more sample efficiency ablations, please refer to the Supplementary Materials.
@@
Parameter Sensitivity. We demonstrate how the key parameters of the BridgePolicy would influence its performance. Specifically, we evaluate different weights $\alpha$ in the overall training loss (8) and the results are shown in Table 4. The results suggest that the distribution aligner can be trained better under a relatively small weight in policy learning. For more ablation study results on the hyperparameters, please refer to the Supplementary Materials.
@@
Modality Fusion. We additionally explore the different ways of fusing the different modalities of the observations. Modality fusion plays a crucial role in resolving the conflict between the multi-modal nature of the observations in policy learning and the one-to-one distribution mapping of the diffusion bridge. We explore the impact of two modality fusion methods, Cross-Attention [lin2022cat] and simple concatenation, on the performance. The results are shown Table 3. When switching from Cross-Attention to simple concatenation, the performance consistently drops on all three evaluated tasks, which demonstrates the effectiveness of using Cross-Attention.
5 Conclusion
@@
In this work, we introduce BridgePolicy, a novel generative policy learning paradigm that directly embeds the observations in the diffusion SDE trajectory via a diffusion bridge formulation rather than merely treating them as high-level conditions in the denoising network like conventional generative policies. This observation-informed SDE enables the sampling to start from a rich and meaningful prior, allowing the policy to more effectively exploit observations, leading to more precise and reliable control. Extensive experiments on simulation and real-world tasks demonstrate its superiority over the current generative policy. The design choices for the distribution aligner and the way of modality fusion, while effective in our experiments, represent one point in a broader design space. We see great potential in further refining these components, and future work will focus on exploring alternative architectures for even more robust distribution alignment and seamless modality integration. In the end, we hope our method helps pave the way towards further research and applications of the diffusion bridge and visuomotor policy learning.
@@ Acknowledgment
Sample from What You See: Visuomotor Policy Learning via Diffusion Bridge with Observation-Embedded Stochastic Differential Equation
Supplementary Materials
@@ Appendix A Overview
We provide additional information, results, and visualization to supplement the main paper. The supplementary materials are organized as follows:
@@
Appendix B Information of experimental setup and real-world data collection
B.1 Real-World Experiment Setup

### v2

@@ 3.2 Modality Fusion and Alignment
Encoding the Robot State and Point Cloud. Our encoding pipeline begins by converting depth images into 3D point clouds, chosen for the efficiency. We then downsample them (512 or 1024 points in simulation; 2048 in real-world) using farthest point sampling (Qi et al., 2017). Finally, a lightweight MLP with LayerNorm and an MLP state encoder map the point cloud and robot state into a shared latent representation for fusion, which can be formulated as:
@@
Multi-Modality Fusion. To enable the diffusion bridge to map between heterogeneous modalities effectively, we introduce the multi-modality fusion module that integrates the 3D point cloud and robot state into a unified representation. This allows the policy to process diverse inputs coherently within our framework. Specifically, we employ the Cross-Attention mechanism (Rombach et al., 2022) to fuse the modalities:
$\boldsymbol{a}_{T}:=\boldsymbol{z}_{obs}=\mathbf{softmax}(\frac{\boldsymbol{z}_{pc}\boldsymbol{z}_{s}^{T}}{\sqrt{d_{s}}})\boldsymbol{z}_{s},$ (6)
@@
We recognize that the terminal state $\boldsymbol{a}_{T}$ of diffusion bridge is an observation representation $\boldsymbol{z}_{obs}$ derived from the learned MLP and fusion via Eq. (5) and Eq. (6), which might introduce errors in the generated actions $\tilde{\boldsymbol{a}}_{0}$ if there exists errors during the encoder mapping. However, we demonstrate that the deviation in the generated action would not be too large, which leads to the following theorem:
Theorem 3.1.
Given the initial value $\boldsymbol{a}_{T}$ and Assumption B.1 on data prediction network $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$. Suppose $\tilde{\boldsymbol{a}}_{T}$ is the perturbed initial value from $\boldsymbol{a}_{T}$, then we have
$\|\tilde{\boldsymbol{a}}_{0}-\boldsymbol{a}_{0}\|\leq C\|\tilde{\boldsymbol{a}}_{T}-\boldsymbol{a}_{T}\|$ (7)
where $C>0$ is a constant, $\boldsymbol{a}_{0}$ and $\tilde{\boldsymbol{a}}_{0}$ are respectively generated from $\boldsymbol{a}_{T}$ and $\tilde{\boldsymbol{a}}_{T}$ via the updating rule Eq. (4).
Please refer to Appendix B for detailed proof. Theorem 3.1 indicates that the error of the action output can be linearly bounded by the error from the observation MLP. Since $C$ is difficult to determine analytically, we provide experimental results to empirically estimate the magnitude of $C$ in Appendix B. In our experiments, the magnitude of $C$ is found to be small, on the order of $10^{-2}$ to $10^{-3}$. Therefore, even when the observation MLP incurs fitting errors, the actions generated from the sampling algorithms of diffusion bridge would be constrained within a narrow range, preventing significant deviation.
Modality Alignment. Although multi-modality fusion aligns the observation and action distributions in shape, significant distributional differences still remain and the learned observation representation lacks sampling capability. To address this, we propose using the contrastive learning loss to train the aligner. The contrastive loss aligns the semantic proximity of the observation and action distributions. Specifically, we adopt CLIP loss (Radford et al., 2021) as the contrastive learning loss:
@@
DP 0.79±0.27 0.31±0.24 0.10±0.12 0.26±0.25 0.45±0.08 0.31±0.17 0.37
DP3 0.87±0.20 0.61±0.29 0.40±0.37 0.51±0.33 0.57±0.08 0.68±0.06 0.60
Simple DP3 0.86±0.12 0.59±0.22 0.38±0.26 0.47±0.28 0.48±0.07 0.68±0.05 0.58
FlowPolicy 0.86±0.16 0.67±0.21 0.59±0.30 0.76±0.26 0.54±0.09 0.70±0.12 0.68
VITA 0.85±0.21 0.58±0.32 0.48±0.48 0.62±0.21 0.55±0.11 0.77±0.19 0.64
BridgePolicy (Ours) 0.91±0.24 0.75±0.28 0.58±0.31 0.79±0.32 0.60±0.07 0.81±0.12 0.74
@@ 4.1 Simulation Experiment
Method Oven-Closing Oven-Opening Pick Place Pour Unplug Average
@@
Simulation Benchmark. In simulation tasks, we evaluate our BridgePolicy on 52 tasks across three benchmarks including Adroit (Rajeswaran et al., 2017), DexArt (Bao et al., 2023), and MetaWorld (Yu et al., 2020). Adroit and DexArt focus on dexterous hand manipulation with varying complexity, while MetaWorld offers a broad spectrum of robotic arm tasks across different difficulty levels.
@@
Expert Demonstration Collection. We collect expert data with script policy in MetaWorld and adopt the VRL3 (Wang et al., 2022) and PPO (Schulman et al., 2017), two reinforcement learning (RL) algorithms, to collect expert data for Adroit and DexArt, respectively. We collect 10 episodes for tasks of the Adroit and the MetaWorld benchmarks and 100 episodes for the DexArt benchmark to train the policies.
@@
Baselines. For comparison, we select state-of-the-art 2D-based methods DP (Chi et al., 2023) which takes the 2D image as its visual input and 3D conditional diffusion or flow based approaches including DP3 (Ze et al., 2024), simple DP3 (its lightweight version), FlowPolicy (Zhang et al., 2025) and VITA (Gao et al., 2025), which generates the action from random Gaussian noise and leverage 3D point cloud as the visual condition in the neural network, as baselines. We set the number of function evaluation (NFE) to 10 for all baselines except FlowPolicy. Based on the settings in the original FlowPolicy paper, we set its NFE to 1 instead of 10 to avoid the error accumulation issue that occurs in multi-step sampling of consistency flow matching (Sabour et al., 2025). For further details of the NFE setting, please refer to Appendix E.2.
@@
Evaluation and Implementation. We run 3 random seeds for each experiment, followed by DP3 (Ze et al., 2024) and FlowPolicy (Zhang et al., 2025). For each random seed, we train the models for 3000 epochs and evaluate the success rate on 20 episodes every 200 training epochs. We pick the five highest success rates per seed and report the mean success rate of the 3 random seeds. For fair comparison, we keep the model architecture consistent with the baselines, the amount of parameters comparable, and the shared hyperparameters, such as learning rates, weight decays of the optimizer, batch sizes, and training epochs. For each task, BridgePolicy is trained on a single NVIDIA RTX 4090 GPU. Please refer to Appendix D for further details of the implementation of BridgePolicy and other baselines.
@@
Performance in Simulator. For Adroit and DexArt, we report the average success rates directly. Particularly for MetaWorld tasks, it is further categorized into four difficulty levels, with average rates reported for each. The quantitative results are reported in Table 1. Our proposed BridgePolicy demonstrates a consistent performance across all evaluated benchmarks, achieving the highest average success rates. All baseline methods show considerably lower success rates, particularly in more challenging settings (Hard and Very Hard in MetaWorld), which also showcases the robustness of our method in tasks with varying complexity. These results validate the effectiveness of BridgePolicy in leveraging the diffusion bridge formulation for improved policy learning in simulation environments. VITA’s performance lies between DP3 and FlowPolicy, potentially because flow matching is performed in a joint image–action latent space while executable actions are obtained with a decoder, which could inevitably introduce generalization error. Refer to Appendix F.4 for detailed success rates of individual tasks.
@@ 4.2 Real-world Experiment
Experiment Setup and Evaluation. We evaluate the BridgePolicy on 5 real-world tasks using the Franka Emika Panda robot. The point cloud is acquired using the ZED-2i camera. We evaluate the policies on 10 episodes of each task. The implementation remains the same as that in the experiments in simulation. We only increase the number of points of the point cloud to 2048 for a more dense representation of the real-world scenario. Real-world experiment environment setup details are provided in Appendix C.
Expert Demonstration Collection. The expert demonstrations are collected by the GELLO human teleopertation system (Wu et al., 2024a), manipulated by an experienced graduate. For each task, we collect 50 episodes to train the policy. Similarly, in the simulator and for a fair comparison, the expert demonstrations provided to all baseline methods and our BridgePolicy are identical. Refer to Appendix C for details of the real-world data collection.
@@
Qualitive Comparison on Real-World Task. Three real-world examples of Oven-Opening and Unplug generated by DP3, FlowPolicy, and BridgePolicy are shown in in Figure 3 and  11. The sequence presents the key frames of Franka finishing the task. The Oven-Opening task consists of 2 key stages: 1) grabbing the oven handle and opening it halfway; 2) moving the arm upon the half-opened door and pushing it down until it is fully opened. In the case of Oven-Opening, DP3 fails to finish the task while FlowPolicy and BridgePolicy succeed. BridgePolicy exactly finishes the 2 stages and fully opens the oven, while FlowPolicy fails to half-open the door in the first stage, and it presses against the handle directly, leaving the door not fully opened. We also observe similar results in the Unplug task, shown in Figure 11. DP3 and FlowPolicy extend their grippers towards the target location, but they do not precisely clamp the socket, thus fail to unplug successfully. BridgePolicy, on the other hand, successfully clamped the plug and pulled it out. Videos of real-world experiments are in Supplementary Materials.
@@
0.0 0.5 1.0 2.0 5.0
Adroit Door 0.79 0.83 0.84 0.84 0.80
DexArt Laptop 0.85 0.86 0.91 0.86 0.89
MW Coffee-Push 0.91 0.99 0.91 0.91 0.92
MW Pick-Place-Wall 0.88 0.93 0.96 0.87 0.83
@@ 4.3 Ablation Study
In this section, we analyze the performance of BridgePolicy under different hyperparameter settings. Additional ablation studies are provided in the Appendix F.
@@
Sample Efficiency. The success rate of the agent in accomplishing tasks largely depends on the number of expert demonstrations. Here, we evaluate how the number of expert demonstrations would affect the success rate of the agent finishing the tasks. As shown in Figure 4, we select four tasks for this ablation study: the MetaWorld Hand-Pull-Side, Pick Place, Stick Pull and the Adroit Pen task. For all four tasks, increasing the number of expert demonstrations generally improves the agent’s success rate. BridgePolicy shows the most stable and robust performance across all demonstration counts among the evaluated tasks, especially with fewer training samples. More sample efficiency ablations can be found in Appendix Section F.
@@
Parameter Sensitivity. We demonstrate how the key parameters of the BridgePolicy would influence its performance. Specifically, we evaluate different weights $\alpha$ in the overall training loss (9) and the results are shown in Table 4. The results suggest that the distribution aligner can be trained better under a relatively small weight in policy learning. More ablation study results on the hyperparameters are in Appendix Section F.
@@
Modality Fusion. We additionally explore the different ways of fusing the different modalities of the observations. Modality fusion plays a crucial role in resolving the conflict between the multi-modal nature of the observations in policy learning and the one-to-one distribution mapping of the diffusion bridge. We explore the impact of two modality fusion methods, Cross-Attention (Lin et al., 2022) and simple concatenation, on the performance. The results are shown in Table 3. When switching from Cross-Attention to simple concatenation, the performance consistently drops on all three evaluated tasks, which demonstrates the effectiveness of using Cross-Attention.
Effectiveness of Diffusion Bridge. We evaluate the effectiveness of Diffusion Bridge formulation compared to standard Diffusion and Regression. We remain the aligner and the modality fusion unchanged and use different policy heads to output actions. The results are shown in Table 5. Regression fails to behave effectively, often collapsing to near-zero success. Standard Diffusion significantly improves the performance over Regression. Diffusion Bridge gain the highest success rate, suggesting that its formulation provides a better-conditioned generation process and yields more precise actions.
Table 5: Quantitative results (Success Rate) to evaluate the effectiveness of Diffusion Bridge in BridgePolicy.
Task Policy Head
Regression Diffusion Bridge
Adroit Pen 0.33 0.43 0.58
Adroit Door 0.00 0.68 0.84
DexArt Laptop 0.16 0.86 0.88
DexArt Faucet 0.00 0.37 0.39
MW Basketball 0.00 0.96 1.00
MW Coffee Pull 0.00 0.86 0.97
5 Conclusion and Limitation
@@
In this work, we present BridgePolicy, a novel generative policy learning paradigm that directly embeds the observations in the diffusion SDE trajectory via a diffusion bridge formulation rather than merely treating them as high-level conditions in the denoising network like conventional generative policies. This observation-informed SDE enables the sampling to start from a rich and meaningful prior, allowing the policy to more effectively exploit observations, leading to more precise and reliable control. BridgePolicy not only outperform the SOTA generative policies, but also extend the application of diffusion bridge to multi-modal and heterogeneous distribution matching. Extensive experiments on simulation and real-world tasks demonstrate its superiority over the current generative policy. Though BridgePolicy outperforms the SOTA generative policies, its sampling remains constrained by its SDE formulation, which limits the possibility of applying distillation methods that would enable one‑step generation.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Overview
We provide additional information, results, and visualization to supplement the main paper. This Appendix are organized as follows:
  • The proof of Theorem 3.1 and its empirical evaluation;
@@
Appendix B Proof of Theorem 3.1 and the Empirical Evaluation
To prove Theorem 3.1, we make the following common assumption about the Lipschitz property of data prediction neural network $\boldsymbol{a}_{\theta}$ (Lu et al., 2022, 2025):
Assumption B.1.
The function $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$ is Lipschitz w.r.t. to its parameters $\boldsymbol{a}_{t}$ and $\boldsymbol{a}_{T}$, i.e., there exists $L_{1},L_{2}>0$ s.t. the following inequality holds,
$\|\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t}^{\prime},\boldsymbol{a}_{T}^{\prime\prime},t)-\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)\|\leq L_{1}\|\boldsymbol{a}_{t}^{\prime}-\boldsymbol{a}_{t}\|+L_{2}\|\boldsymbol{a}_{T}^{\prime\prime}-\boldsymbol{a}_{T}\|.$ (10)
Theorem 3.1. Given the initial value $\boldsymbol{a}_{T}$ and Assumption B.1 on data prediction network $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$. Suppose $\tilde{\boldsymbol{a}}_{T}$ is the perturbed initial value from $\boldsymbol{a}_{T}$, then we have
$\|\tilde{\boldsymbol{a}}_{0}-\boldsymbol{a}_{0}\|\leq C\|\tilde{\boldsymbol{a}}_{T}-\boldsymbol{a}_{T}\|$ (7)
where $C>0$ is a constant, $\boldsymbol{a}_{0}$ and $\tilde{\boldsymbol{a}}_{0}$ are respectively generated from $\boldsymbol{a}_{T}$ and $\tilde{\boldsymbol{a}}_{T}$ via the updating rule Eq. (4).
Proof.
Consider the two updating rules from time step $s$ to $t\in[0,s]$:
$\begin{gathered}\boldsymbol{a}_{t}=c_{1,t:s}\boldsymbol{a}_{s}+c_{2,t:s}\boldsymbol{a}_{T}+c_{3,t:s}\boldsymbol{a}_{\theta}(\boldsymbol{a}_{s},\boldsymbol{a}_{T},s)+\delta^{d}_{t:s}\boldsymbol{\epsilon},\\ \tilde{\boldsymbol{a}}_{t}=c_{1,t:s}\tilde{\boldsymbol{a}}_{s}+c_{2,t:s}\tilde{\boldsymbol{a}}_{T}+c_{3,t:s}\boldsymbol{a}_{\theta}(\tilde{\boldsymbol{a}}_{s},\tilde{\boldsymbol{a}}_{T},s)+\delta^{d}_{t:s}\boldsymbol{\epsilon},\end{gathered}$ (11)
where we simply denote the three coefficients w.r.t. $\boldsymbol{a}_{s}$, $\boldsymbol{a}_{T}$, and $\boldsymbol{a}_{\theta}$ as $c_{1,t:s}$, $c_{2,t:s}$, and $c_{3,t:s}$, respectively. Define the error at time $t$ as
$\boldsymbol{e}_{t}\triangleq\|\tilde{\boldsymbol{a}}_{t}-\boldsymbol{a}_{t}\|.$ (12)
Then we have
$\boldsymbol{e}_{t}$ $=\|\tilde{\boldsymbol{a}}_{t}-\boldsymbol{a}_{t}\|$ (13)
$=\|c_{1,t:s}(\tilde{\boldsymbol{a}}_{s}-\boldsymbol{a}_{s})+c_{2,t:s}(\tilde{\boldsymbol{a}}_{T}-\boldsymbol{a}_{T})+c_{3,t:s}[\boldsymbol{a}_{\theta}(\tilde{\boldsymbol{a}}_{s},\tilde{\boldsymbol{a}}_{T},s)-\boldsymbol{a}_{\theta}(\boldsymbol{a}_{s},\boldsymbol{a}_{T},s)]\|$
$\leq^{(i)}c_{1,t:s}\|\tilde{\boldsymbol{a}}_{s}-\boldsymbol{a}_{s}\|+c_{2,t:s}\|\tilde{\boldsymbol{a}}_{T}-\boldsymbol{a}_{T}\|+c_{3,t:s}\|\boldsymbol{a}_{\theta}(\tilde{\boldsymbol{a}}_{s},\tilde{\boldsymbol{a}}_{T},s)-\boldsymbol{a}_{\theta}(\boldsymbol{a}_{s},\boldsymbol{a}_{T},s)\|$
$\leq^{(ii)}c_{1,t:s}\boldsymbol{e}_{s}+c_{2,t:s}\boldsymbol{e}_{T}+c_{3,t:s}(L_{1}\boldsymbol{e}_{s}+L_{2}\boldsymbol{e}_{T})$
$=(c_{1,t:s}+c_{3,t:s}L_{1})\boldsymbol{e}_{s}+(c_{2,t:s}+c_{3,t:s}L_{2})\boldsymbol{e}_{T}$
$=C_{1,t:s}\boldsymbol{e}_{s}+C_{2,t:s}\boldsymbol{e}_{T},$
where $(i)$ follows from the triangle inequality, $(ii)$ follows from Eq. (10), and we simply denote $C_{1,t:s}\triangleq c_{1,t:s}+c_{3,t:s}L_{1}$ and $C_{2,t:s}\triangleq c_{2,t:s}+c_{3,t:s}L_{2}$. Given $M+1$ time steps $\left\{t_{i}\right\}_{i=0}^{M}$ decreasing from $t_{0}=T$ to $t_{M}=0$, repeating the argument (13), we have
$\boldsymbol{e}_{0}$ $\leq C_{1,t_{M}:t_{M-1}}\boldsymbol{e}_{t_{M-1}}+C_{2,t_{M}:t_{M-1}}\boldsymbol{e}_{T}$ (14)
$\leq C_{1,t_{M}:t_{M-1}}C_{1,t_{M-1}:t_{M-2}}\boldsymbol{e}_{t_{M-2}}+(C_{1,t_{M}:t_{M-1}}C_{2,t_{M-1}:t_{M-2}}+C_{2,t_{M}:t_{M-1}})\boldsymbol{e}_{T}$
$\leq\cdots$
$\leq\prod_{i=1}^{M}C_{1,t_{M-i+1}:t_{M-i}}e_{t_{0}}+\sum_{i=1}^{M}C_{2,t_{M-i+1}:t_{M-i}}\prod_{k=i+1}^{M}C_{1,t_{M-k+1}:t_{M-k}}\boldsymbol{e}_{T}$
$=\left(\prod_{i=1}^{M}C_{1,t_{M-i+1}:t_{M-i}}+\sum_{i=1}^{M}C_{2,t_{M-i+1}:t_{M-i}}\prod_{k=i+1}^{M}C_{1,t_{M-k+1}:t_{M-k}}\right)\boldsymbol{e}_{T}.$
Hence,
$\quad\boldsymbol{e}_{0}\leq C\boldsymbol{e}_{T}$ (15)
$\Leftrightarrow$ $\quad\|\tilde{\boldsymbol{a}}_{0}-\boldsymbol{a}_{0}\|\leq C\|\tilde{\boldsymbol{a}}_{T}-\boldsymbol{a}_{T}\|,$
where $C\triangleq\prod_{i=1}^{M}C_{1,t_{M-i+1}:t_{M-i}}+\sum_{i=1}^{M}C_{2,t_{M-i+1}:t_{M-i}}\prod_{k=i+1}^{M}C_{1,t_{M-k+1}:t_{M-k}}$ is a constant only depending on the finite number of sampling steps and Lipschitz constants in Assumption B.1, which concludes the proof.
Empirical Estimation of Constant Parameter $C$ in Theorem 3.1. We conduct experiments on six tasks to empirically evaluate the magnitude of $C$. Specifically, we add different level of random perturbations to the same learned observation representation $\boldsymbol{a}_{T}=\boldsymbol{z}_{obs}$ and generate actions through the reverse process, after which we calculate the corresponding $C$. Here $M=10$ is consistent with NFE in the simulation experiments. For each task, we average the results on 1000 sampling trials.
Figure 5: Empirical Evaluation of Constant $C$ in Theorem 3.1.
According to Theorem 3.1, after taking the logarithm of both sides of the inequality ($\log\|\tilde{\boldsymbol{a}}_{0}-\boldsymbol{a}_{0}\|\leq\log C+\log\|\tilde{\boldsymbol{a}}_{T}-\boldsymbol{a}_{T}\|$), the corresponding error curve should lie below a straight line with a slope of approximately $1$. Our experimental results are consistent with the statement of Theorem 3.1. In Figure 5, all fitted lines exhibit slopes roughly in the range of 1.091.231.09–1.23 and the constant $C$ falls in $10^{-2}$ and $10^{-3}$, which aligns with our formulation. The errors introduced by the observation MLP would not influence the generated action $\boldsymbol{a}_{0}$ too much.
@@
Appendix C Experimental Setup and Real-World Data Collection Details
C.1 The Simulation and Real-World Task Example
The examples of the simulation and real-world tasks are shown in Figure 6.
Figure 6: Examples of the simulation and real-world tasks.
C.2 Real-World Experiment Setup